Find the number of molecules of 3.01x1023 mole of HCI.
1
Expert's answer
2020-04-20T15:07:30-0400

I dare to suggest that 3.01*1023 is not a number of moles but number of molecules, so the task is reverse - to derive number of moles from number of molecules, check it please (too large amount of moles, usually such quantities represent amount of single particles - atoms, molecules, not moles).

Anyway

N = n*Na(Avogadro constant)

  • if n = 3.01*1023 (my assumption is wrong and your task is right) then

N = 3.01*1023*6.02*1023 = 1.812*1047 molecules

  • if N = 3.01*1023 (my assumption is correct) then

n = 3.01*1023 / 6.02*1023 = 0.5 moles